Willow Creek Manufacturing, Inc. issued $7,000,000 face value of bonds at 99 on January 1, 2019. The bonds are dated January 1, 2019, pay interest semi-annually at 6% on June 30 and December 31, and mature in 10 years. Straight-line amortization is used for discounts and premiums. On January 1, 2024, $4,200,000 of the bonds are called at 101.

1. The cash paid by Willow Creek Manufacturing when the bonds are redeemed would be...

2. The gain or loss on the bond redemption recorded by Willow Creek Manufacturing on 1/1/24 would be...
Answer #1
1 Cash paid when redeemed = $4,200,000 x 101/100 = $4,242,000
2 Issue price of bonds = $7,000,000 x 99/100 = $6,930,000
Discount on bonds payable = $7,000,000 - $6,930,000 = $70,000
Discount on bonds payable pertaining to bonds redeemed = $70,000 x $4,200,000/$7,000,000 = $42,000
Discount amortized upto January 1, 2024 = $42,000 x 10/20 = $21,000
Balance in Discount on bonds payable on January 1, 2024 = $42,000 - $21,000 = $21,000
Carrying value of bonds redeemed = $4,200,000 - $21,000 = $4,179,000
Loss on the bond redemption = $4,242,000 - $4,179,000 = $63,000